A sample of 14 randomly selected commuters in Chicago showed the average of the commuting times was 33.2 minutes. If the standard decision was 8.3 minutes, find the confidence interval of the true mean.
Answer #1

As confidence interval not specified, we assume it for 95%

n = 14

mean = 33.2

t-value of 95% CI = 2.1604

std. dev. = 8.3000

SE = std.dev./sqrt(n)

= 8.3/sqrt(14)

= 2.21827

ME = t*SE

= 2.1604*2.21827

= 4.79228

Lower Limit = Mean - ME = 33.2 - 4.79228 = 28.40772

Upper Limit = Mean + ME = 33.2 + 4.79228 = 37.99228

95% CI (28.4077 , 37.9923)